OpenStudy (anonymous):

I do not know AT ALL how to solve this. Can someone work it out, give me the answer, and explain how they got their answer? Simplify: 4 square root of 2 end root plus 7 square root of 2 end root minus 3 square root of 2

OpenStudy (anonymous):

A.)2 square root of 8 B.)8 square root of 2 C.)8 square root of 6 D.) 6 square root of 8

OpenStudy (mathmale):

Could you possibly type this problem in Equation Editor or draw it in the Draw utility? Otherwise many of the rest of us have to try guessing what you mean. Please note: Please do not ask for (and do not expect) answers. Those who tutor on OpenStudy are bound to help you through problems (enabling you to solve them on your own), and are forbidden to give you answers.

OpenStudy (mathmale):



OpenStudy (mathmale):

\[4\sqrt{2}+ ... ?\]

OpenStudy (amistre64):

4x = x+x+x+x 7x = x+x+x+x+x+x+x -3(x) = 3(-x)= -x-x-x since its all in terms of some value x ... how many xs do we end up with?

OpenStudy (amistre64):

and yeah, notation is useful

OpenStudy (mathmale):

4 square root of 2 end root plus 7 square root of 2 end root minus 3 square root of 2\[4\sqrt{2}+7\sqrt{2}-3\sqrt{2}=?\]

OpenStudy (mathmale):

Each term has Sqrt(2) in it. Why not factor out that Sqrt(2)? What's left are three terms to be combined into one: (4+7-3). Please finish this.